# Theory of Super-Resolution Data Assimilation with Conditional Variational Autoencoders: Using Super-Resolution Operators as Background Error Covariance Matrices

### v3

@@ V.4.1 Training method
In principle, no HR simulation is required because the training data include LR forecasts rather than HR results (Sections IV.4 and IV.5). These LR forecasts are obtained by numerical integration in which the analysis states are used as initial conditions. In offline training, however, such an analysis is unavailable because the CVAE is not trained yet. In realistic applications, these initial conditions can be given from existing assimilated data, such as reanalysis,Hersbach et al. (2020); KOSAKA et al. (2024) or online training will be effective,Rasp (2020) where neural-network training and data generation are executed alternately.
@@
The SR model was trained, followed by training the CVAE. During the CVAE training, all parameters in the SR model were fixed, as discussed in Section IV.5. The loss functions of the SR model and the CVAEs are $l_{\text{SR}}$ in Eq. (18) and $l_{\text{SRDA}}$ in Eq. (13), respectively. The CVAE and the SR model were trained using the Adam optimizerKingma and Ba (2015) with a learning rate of $1\times 10^{-4}$ and a mini-batch size of 128. Each training was terminated using early stopping with a patience parameter of 50 epochs. The parameters in $l_{\text{SRDA}}$ were set to $r=1.2\times 10^{-5}$ and $b=1.0\times 10^{-3}$. The value of $r$ was given by the variance of the measurement errors (Section V.2), while $b$ was determined through grid search to balance the reconstruction error and the KL divergence in the ELBO. Without this balancing, the performance of the CVAE would deteriorate. This phenomenon is often observed in VAEs,Higgins et al. (2017); Alemi et al. (2018) and the loss balancing can be optimized during training.Asperti and Trentin (2020) For simplicity, however, such an advanced technique was not employed in the present study.
V.4.2 Testing method
@@
The sequential SRDA was tested using the method described in Section IV.6. Figure 6(b) shows a schematic of the test phase. The only difference between Figs. 4 and 6(b) is that observations are generated by subsampling the HR simulation results. Due to this difference, the HR results are regarded as the ground truth, and the SRDA can be evaluated by comparing its inference to the ground truth. We newly conducted 500 HR fluid simulations. For each case, the LR fluid model was initialized with the low-pass filtered HR vorticity at $t=0$; the SRDA was subsequently conducted until $t=20$ with the assimilation interval $\Delta T=1$. Without SRDA, the vorticity evolution would differ from the ground truth, not only due to the difference in the initial condition (i.e., all high-wavenumber components are removed), but also due to the difference in the fluid model resolution (i.e., LR versus HR).
Two test metrics were used to assess pixel-wise accuracy and pattern consistency, both of which measure deviations from the ground truth. The mean absolute error (MAE) ratio quantifies pixel-wise errors.
@@
where the summation is taken over all grid points, $\omega_{i}$ denotes the ground-truth vorticity at the $i$-th grid point, and $\hat{\omega}_{i}$ is the corresponding inference. The mean structural similarity index measure (MSSIM) loss evaluates the consistency of spatial patterns:Wang et al. (2004)
@@
where ${\text{C}_{1}}=0.01$, ${\text{C}_{2}}=0.03$, $\mu_{i}$ and $\sigma_{i}^{2}$ are the mean and variance of the ground truth, respectively, $\hat{\mu}_{i}$ and $\hat{\sigma}_{i}^{2}$ are the corresponding quantities for the inference, and $\gamma_{i}$ is the covariance between the ground truth and the inference. The variables in the summation are calculated locally in space using a Gaussian filter. The MSSIM loss takes a value greater than or equal to 0; a smaller value indicates spatial patterns more similar to the ground truth. A detailed discussion of MSSIM can be found in Wang et al.Wang et al. (2004)
@@
V.5 Baseline model using Ensemble Kalman Filter (EnKF)
As a comparison to the SRDA, we employed an ensemble Kalman filter (EnKF)Evensen (1994) for the LR fluid model, using the perturbed observation method.Burgers, van Leeuwen, and Evensen (1998) The choice of EnKF was motivated by its ability to estimate uncertainty through ensemble simulations. The uncertainty with the EnKF is compared with that with the SRDA.
Assimilation was conducted in the HR space after the LR forecast was super-resolved to the HR using bicubic interpolation, where the assimilation interval $\Delta T=1$ is the same as in the SRDA. Barthélémy et al.Barthélémy et al. (2022) argued that performing the EnKF in the HR space can yield more accurate inferences than those in the LR space when observations are defined at HR grid points. Furthermore, this approach clarifies the comparison with the SRDA because both EnKF and SRDA employ the LR fluid model and perform DA in the HR space.
The background error covariance matrices were spatially localized using a function proposed by Gaspari and Cohn.Gaspari and Cohn (1999) These error covariances were inflated by adding Gaussian noise to the analysis, prior to the forecast process.Whitaker et al. (2008) This Gaussian noise incorporates spatial correlation by estimating its covariance from the training data at each time step. The initial perturbations for creating ensemble members were also drawn from this Gaussian distribution. We confirmed that the results are not sensitive to the statistics of the Gaussian noise for inflation. Indeed, similar results were obtained when the Gaussian noise had a constant correlation length.
Hyperparameters for the EnKF were adjusted using the training data to minimize the MAE. Specifically, the following parameters were tuned: the number of ensemble members (set to 300 after tuning), the amplitude of the initial perturbation for generating ensemble members, the amplitude of the additive inflation, the localization radius, and the amplitude for perturbing observations. The full implementation of the EnKF is also available at our GitHub repository (see Data Availability Statement).
VI Results and discussion
VI.1 Time evolution of vorticity in the jet system
Before discussing the SRDA results, we describe the time evolution of HR vorticity in the jet system and confirm the sensitivity to initial conditions (i.e., a positive Lyapunov exponent).
Figure 7 shows a typical evolution of HR vorticity. Initially ($t=0$), an unstable jet exists around the $y$ center with small perturbations. The jet starts meandering and breaks down into multiple vortices, accompanied by the development of fine filaments ($1\lesssim t\lesssim 4$). Subsequently, vortex merging occurs, intensifying small-scale structures again ($4\lesssim t\lesssim 12$). Finally, a wavenumber-1 structure continues to propagate westward ($12\lesssim t$). This final state is statistically steady and can be understood by Rossby wave dynamics.David, Marshall, and Zanna (2017)
@@ VI.1 Time evolution of vorticity in the jet system
We confirmed that the maximum Lyapunov exponent $\lambda$ is positive: $\lambda=3.11$. This suggests that the temporal evolution of vorticity exhibits initial-value sensitivity, likely due to the barotropic instability of the initial conditions. Indeed, without DA, the deviation from the ground truth continues to increase over time, while the SRDA effectively suppresses this error growth (Section VI.3).
VI.2 Accuracy of SR
We first confirm the accuracy of inference by the SR model, as the SRDA is based on this model. Figure 8 shows three sets of snapshots for the ground truth, observations, LR input, and SR output. The results indicate that the SR model reproduces the fine filaments and small-scale structures within vortices. These patterns are not evident in the LR inputs or observations.
@@ VI.2 Accuracy of SR
Flow-field reconstruction has been extensively studied using neural networks.Lluís et al. (2020); Maulik et al. (2020); Gundersen et al. (2021); Wang et al. (2022) In these studies, inputs are incomplete, such as spatially sparse sensor measurements, where only a portion of the entire domain is available. Neural networks are then trained by comparing their output to complete target data. The setup of our problem is opposite to that of these studies: in our case, the input is complete, while the target (i.e., the observation) is incomplete.
@@
The success of the SR in our study is attributable to two factors: (i) the point-wise nature of the loss function and (ii) the weight sharing of convolution. The loss function $l_{\text{SR}}$ in Eq. (18) operates at individual grid points, and only the non-missing observed points contribute to $l_{\text{SR}}$. This property enables backpropagation even when the target data are missing, as discussed in Section IV.4. Furthermore, since convolution shares its kernel across the space (known as weight sharing),Aloysius and Geetha (2017) the efficient learning was realized using the small amount of non-missing observational data.
@@
Although the inference appears to be accurate, we confirmed that the SR model cannot adequately correct forecast errors. One might think that the SR output could be used as the analysis due to its high accuracy (Fig. 8). However, we found that test errors accumulated over time when the downsampled SR output was used as the initial condition for the LR fluid model. This result indicates that observations need to be assimilated into forecasts.
VI.3 Accuracy of LR forecast by SRDA
@@
We demonstrate that the sequential SRDA can make more accurate inferences than the EnKF. The LR forecast is compared in this section, and the HR analysis is examined next (Section VI.4).
@@
Figure 9 shows snapshots of the ground truth and forecasts for the EnKF and the SRDA. To align the resolution, a low-pass filter was applied to the HR ground-truth vorticity. The figure also includes LR simulation results without SR or DA. During the jet collapse ($t=2$), differences in the inferences are not obvious, likely because the LR fluid model cannot well emulate fine filaments. During the vortex merging ($t=9$), the SRDA most accurately describes the vortex-split distribution near the center. The EnKF also captures this pattern, but the SRDA inference is closer to the ground truth. Without SR or DA, the vortex-split structure is not reproduced. In the steady state ($t=16$), both EnKF and SRDA accurately infer the vorticity phase. Without SR or DA, the vortex shapes and locations still differ from the ground truth.
@@
Figure 9: Vorticity snapshots of the ground truth and forecasts for the EnKF and the SRDA at $t=2$, 9, and 16. For comparison, the figure also shows LR simulation results without SR or DA. A low-pass filter was applied to the ground-truth snapshots to align the resolution to the LR.
@@
The quantitative analysis further confirms the accuracy of the SRDA forecast. The top panels in Fig. 10 show the time series of the mean values for the MAE ratio and the MSSIM loss. Without SR or DA (green dotted lines), the test errors increase over time due to the sensitivity to initial conditions (i.e., the positive Lyapunov exponent in Section VI.1). In contrast, this error growth is suppressed by the EnKF (orange dashed lines) and the SRDA (blue solid lines). The error reduction occurs every time unit since the assimilation interval was set to 1. When averaged over time, both time series have two peaks, corresponding to the two developments of small-scale patterns by the jet collapse ($t\sim 2$) and the vortex merging ($t\sim 9$). The bottom panels in Fig. 10 show the differences in the test errors between the SRDA and the EnKF, where the gray areas represent the 10th to 90th percentile values. The test errors of the SRDA are significantly smaller than those of the EnKF. Thus, we can conclude that the SRDA forecast outperforms the EnKF in terms of pixel-wise accuracy and pattern consistency.
@@
Figure 10: Time series of the MAE ratio and the MSSIM loss for the forecast states in the 500 test simulations. The top panels show the mean test errors at each time step with an output interval of 0.25. The bottom panels show the differences between the SRDA and the EnKF, where the gray areas cover the 10th to 90th percentile values and the solid lines indicate the mean differences.
@@
Within the SRDA framework, the LR forecast can be super-resolved to the HR at an arbitrary time by the SR model (Section IV.6). Figure 11 shows a typical vortex merging between $t=8.5$ and 11.0 from the ground truth, the LR forecast, and the super-resolved forecast. The merging of two vortices is not obvious in the LR, but is well captured in the super-resolved forecast.
@@
Figure 11: Typical vortex merging in the ground truth, the LR forecasts, and the super-resolved forecast.
VI.4 Accuracy of HR analysis by SRDA
@@
We next demonstrate the accuracy of the HR analysis provided by the SRDA. Figure 12 shows snapshots of the ground truth and analyses. During the jet collapse ($t=2$), the filament structure appears much more distinct in the SRDA analysis than in the EnKF. Furthermore, during the vortex merging ($t=9$ and 11), the SRDA snapshots exhibit sharper vortex shapes and clearer internal structures compared to the EnKF.
@@
Figure 12: Vorticity snapshots of the ground truth and analyses for the EnKF and the SRDA at $t=2$, 9, and 11.
@@
The quantitative analysis also confirms the accuracy of the SRDA analysis. Figure 13 shows the time series of the MAE ratio and the MSSIM loss, as in Fig. 10. The dots represent the assimilation time steps, with assimilation intervals at 1. The SRDA and EnKF analyses are inferred only at these times. The EnKF time series (orange dashed lines) show two peaks, corresponding to the two developments of small-scale patterns linked to the jet collapse ($t\sim 2$) and the vortex merging ($t\sim 9$). These peaks are less apparent in the SRDA time series (blue solid lines), especially for the first peak. This result suggests that the SRDA effectively mitigates the error increase during the jet collapse. The bottom panels in Fig. 13 indicate that the MAE ratio and the MSSIM loss for the SRDA are significantly smaller than those of the EnKF. Thus, we can conclude that the SRDA analysis is more accurate than the EnKF analysis, in line with the forecast results (Section VI.3).
@@
Figure 13: Time series of the MAE ratio and the MSSIM loss for the analysis states in the 500 test simulations. These states were inferred only at the analysis time steps, denoted by dots. The top panels show the mean test errors. The bottom panels show the differences between the SRDA and the EnKF, where the gray areas cover the 10th to 90th percentile values and the solid lines indicate the mean differences.
@@
We compare uncertainties estimated by the EnKF and the SRDA. Here, standard deviation (SD) is used as an indicator of uncertainty. For the SRDA, SD was directly calculated by taking the square root of the estimated variances in Eq. (12), while for the EnKF, SD at each grid point was calculated from all ensemble members (i.e., 300 members).
@@
Figure 14 compares distributions of the analysis values and their SDs for the EnKF and the SRDA. The SRDA SD tends to be large around vortex filaments or vortex edges. This tendency appears to be most evident in the jet collapse ($t=2$). The EnKF SD also exhibits a similar tendency; however, the SD distribution is blurred, likely reflecting the blurred vortex patterns. The SD value ranges also differ: the EnKF SD is between 0.3 and 0.9, while the SRDA SD is between 0.8 and 0.9. Related to this range difference, the SRDA tends to overestimate SD in regions where vorticity is close to zero, which can be found at $t=9$ and 11 in Fig. 14. In contrast, the EnKF SD tends to be small in small-vorticity regions. This overestimation of the SRDA may be due to a simplified variance estimation in the present CVAE (Section V.3), where the variances are computed using the SR model from the LR analysis. Therefore, although the SRDA demonstrated the potential for uncertainty quantification without relying on ensemble calculation, further model development is necessary. Recent studies demonstrate that VAEs can learn error distributions of model forecasts,Grooms (2021); Yang and Grooms (2021); Grooms et al. (2023); Xiao et al. (2024) and their model architectures may be effective for the SRDA.
@@
Figure 14: Vorticity snapshots of the analyses and their standard deviations (SDs) for the EnKF and the SRDA at $t=2$, 9 and 11. The analyses shown here are the same as in Fig. 12. To compare them with their SDs, these panels are displayed again.
VI.5 Computation time of SRDA

### v4

@@ 5.4.1 Training method
In principle, HR simulations are not required because the training data include LR forecasts rather than HR results. These LR forecasts are obtained by numerical integration in which the analysis states are used as initial conditions. In offline training, however, such an analysis is unavailable because the CVAE is not trained yet. Thus, the filtered HR results were used as the initial conditions. In realistic applications, these initial conditions can be obtained from existing assimilated data [61, 62], or online training would be effective [63, 64], where training and data generation are performed alternately or in parallel.
@@
The SR model was trained, followed by training the CVAE. For both trainings, the Adam optimizer [65] was used with a learning rate of $1\times 10^{-4}$ and a mini-batch size of 128. Each training was terminated using early stopping with a patience parameter of 50 epochs. During the CVAE training, we fixed all parameters in the SR model. The loss functions of the SR model and the CVAEs are $l_{\text{SR}}$ in Eq. (18) and $l_{\text{SRDA}}$ in Eq. (13), respectively. The parameters in $l_{\text{SRDA}}$ were set to $r=1.2\times 10^{-5}$ and $b=1.0\times 10^{-3}$. The value of $r$ was determined by the variance of the measurement errors, while $b$ was determined through grid search to balance the reconstruction error and the KL divergence in the ELBO. Without this balancing, the performance of the CVAE would deteriorate. This phenomenon frequently occurs in VAEs [66, 67], and the loss balancing can be optimized during training [68]. For simplicity, however, this advanced technique was not employed in our study.
5.4.2 Testing method
@@
The sequential SRDA was tested using the method described in Section 4.6. Figure 8b shows a schematic of the test phase. The only difference between Figs. 6 and 8b is that observations are generated by subsampling the HR simulation results. Due to this difference, the HR results are regarded as the ground truth, and the SRDA is evaluated by comparing its inferences to the ground truth. We newly conducted 500 HR fluid simulations. For each case, the LR fluid model was initialized with the low-pass filtered HR vorticity at $t=0$; the SRDA was subsequently conducted until $t=20$ with the assimilation interval $\Delta T=1$. Without SRDA, the vorticity evolution would significantly differ from the ground truth, not only due to differences in the initial condition (i.e., all high-wavenumber components are completely removed) but also due to the difference in the fluid model resolution (i.e., LR versus HR).
Two test metrics were used to assess pixel-wise accuracy and pattern consistency, both of which measure deviations from the ground truth. These two types of metrics are frequently used in SR studies [29] because they complement each other. The pixel-wise accuracy quantifies localized differences but may overestimate discrepancies when spatial patterns are correctly inferred with a slight shift. The pattern consistency mitigates this effect by evaluating structural similarity between inferred and ground-truth patterns, although it may overlook significant pixel-wise deviations.
The mean absolute error (MAE) ratio quantifies pixel-wise errors.
@@
where $\omega_{i}$ denotes the ground-truth vorticity at the $i$-th grid point, $\hat{\omega}_{i}$ is the corresponding inference, and the summation is taken over all grid points.
The mean structural similarity index measure (MSSIM) loss evaluates the consistency of spatial patterns [69]:
@@
where ${\text{C}_{1}}=0.01$, ${\text{C}_{2}}=0.03$, $\mu_{i}$ and $\sigma_{i}^{2}$ are the mean and variance of the ground truth, respectively, $\hat{\mu}_{i}$ and $\hat{\sigma}_{i}^{2}$ are the corresponding quantities for the inference, and $\gamma_{i}$ is the covariance between the ground truth and the inference. These variables in the summation are calculated locally in space using a Gaussian filter. The MSSIM loss takes a value greater than or equal to zero; a smaller value indicates spatial patterns more similar to the ground truth. A detailed discussion of MSSIM can be found in Wang et al. [69].
@@
5.5 Baseline model using Ensemble Kalman Filter (EnKF)
As a baseline for comparison with the SRDA, we employed an ensemble Kalman filter (EnKF) [70] for the LR fluid model, using the perturbed observation method [71]. The choice of EnKF was motivated by its ability to estimate uncertainty through ensemble simulations. The uncertainties estimated by the EnKF and SRDA are compared in Section 6.4.
Assimilation was conducted in the HR space after the LR forecast was super-resolved to the HR using bicubic interpolation, with the assimilation interval $\Delta T$ ($=1$) being the same as in the SRDA. Barthélémy et al. [34] argued that performing the EnKF in the HR space can yield more accurate inferences than those in the LR space when observations are defined at HR grid points. Furthermore, this approach clarifies the comparison with the SRDA because both EnKF and SRDA employ the LR fluid model and perform DA in the HR space.
The background error covariance matrices were spatially localized using a function proposed by Gaspari and Cohn [72]. These error covariances were inflated by adding Gaussian noise to the analysis before the forecast process [73]. This Gaussian noise incorporates spatial correlation by estimating the covariance from the training data at each time step. The initial perturbations for creating ensemble members were also drawn from this Gaussian distribution.
Hyperparameters for the EnKF were adjusted using the training data to minimize the MAE. Specifically, the following parameters were tuned: the number of ensemble members (set to 300 after tuning), the amplitude of the initial perturbation for generating ensemble members, the amplitude of the additive inflation, the localization radius, and the amplitude for perturbing observations. The full implementation of the EnKF is also available at our Zenodo repository (see Data Availability Statement).
6 Results and discussion
6.1 Time evolution of vorticity in the jet system
We first describe the time evolution of HR vorticity in the jet system and confirm the sensitivity to initial conditions (i.e., a positive Lyapunov exponent).
Figure 9 shows a typical evolution of HR vorticity. Initially ($t=0$), an unstable jet exists around the $y$ center with small perturbations. The jet starts meandering and breaks down into multiple vortices, accompanied by the development of fine filaments ($1\lesssim t\lesssim 4$). Subsequently, vortex merging occurs, intensifying small-scale structures again ($4\lesssim t\lesssim 12$). Finally, a wavenumber-1 structure continues to propagate westward ($12\lesssim t$). This final state is statistically steady and can be understood by Rossby wave dynamics [51].
@@ 6.1 Time evolution of vorticity in the jet system
The maximum Lyapunov exponent $\lambda$ is positive: $\lambda=3.11$. This indicates that the temporal evolution of vorticity exhibits initial-value sensitivity, likely due to the barotropic instability. Without DA, the deviation from the ground truth continues to increase over time, as shown in Section 6.3.
6.2 Accuracy of SR inference
We show the accuracy of inference by the SR model, as the SRDA is based on this model. Figure 10 shows three sets of snapshots for the ground truth, observations, LR input, and SR output. The SR model reproduces the fine filaments and the small-scale structures within vortices. These patterns are not evident in the LR inputs or observations.
@@ 6.2 Accuracy of SR inference
Flow-field reconstruction has been extensively studied using neural networks [74, 75, 76, 77]. In these studies, inputs are incomplete, such as spatially sparse sensor measurements. Neural networks are then trained by comparing their outputs to complete target data. Our problem setup is opposite to that of these studies: in our case, the input is complete, while the target (i.e., the observation) is incomplete.
@@
The success of SR in our study is attributable to two factors: (i) the point-wise nature of the loss function and (ii) the weight sharing of convolution. The loss function $l_{\text{SR}}$ in Eq. (18) operates at individual grid points, and only the non-missing observed points contribute to $l_{\text{SR}}$. This point-wise nature enables backpropagation even when the target data are missing. Furthermore, since convolution shares its kernel across space (known as weight sharing) [78], the efficient learning was realized using the small amount of non-missing observational data.
@@
Although the inference appears to be accurate, the SR model cannot adequately correct forecast errors. One might think that the SR output could be used as the analysis due to its high accuracy (Fig. 10). However, test errors accumulated over time when the downsampled SR output was used as the initial condition for the LR fluid model. This result indicates that observations need to be assimilated into forecasts.
The trained SR model makes non-local inference. This non-locality is essential for assimilating observations, as discussed in Section 4.3. We use two indicators for interpreting the SR model: gradients and local attribution maps (LAMs) [31]. Gradients are expressed as $\partial\mathsf{F}(\bm{x})/\partial\bm{x}$, where $\mathsf{F}$ is the trained SR model. Such a simple gradient is sometimes not a good indicator for interpreting how a model output depends on its input due to nonlinearity [31, 79]. LAM, proposed by Gu and Dong [31], combines gradient methods with path integral techniques for interpreting behaviors of nonlinear SR models.
Figure 11 shows gradient and LAM fields for the three SR samples previously discussed in Fig. 10. We focus on the square regions (black boxes) in the figures. The gradients and LAMs indicate that colored points in the LR inputs contribute to the HR outputs in these boxes. If there are observed values in these boxes, the colored points in the inputs are updated by minimizing the loss function (Section 4.3). Thus, observations have a non-local influence through the trained SR model. This is an essential reason for the success of the proposed SRDA. Interestingly, the LAMs tend to take larger values over wider regions, compared to the gradients (Fig. 11), suggesting the nonlinearity of the SR model. For instance, in the first row of the figure, the LAM has a larger value at one grid point below the black box, matching the position of the vortex filament.
Figure 11: Attribution maps for SR. The left two columns, SR output and LR input, are the same as in Fig. 10, but shown here for comparison. Local attribution maps (LAMs) [31] were calculated for the small square regions of $11\times 1$ 11 (i.e., black boxes in the figure). These boxes were specified on the HR grids, so they are not necessarily aligned with the LR grids. The gradients and LAMs were normalized by their maximum values, with ranges between 0 and 1, following Gu and Dong [31].
6.3 Accuracy of LR forecast
@@
We demonstrate that the sequential SRDA can make more accurate inferences than the EnKF. The LR forecast is discussed here, and the HR analysis is examined next.
@@
Figure 12 shows snapshots of the ground truth and forecasts for the EnKF and SRDA. To match the resolution, a low-pass filter was applied to the HR ground-truth vorticity. The figure also includes LR simulation results without SR or DA. During the jet collapse ($t=2$), differences in the inferences are not obvious, likely because the LR fluid model cannot well emulate fine filaments. During the vortex merging ($t=9$), the SRDA most accurately describes the vortex-split distribution near the center. The EnKF also captures this pattern, but the SRDA inference is closer to the ground truth. Without SR or DA, the vortex-split structure is not reproduced. In the steady state ($t=16$), both EnKF and SRDA accurately infer the vorticity phase. Without SR or DA, the vortex shapes and locations still differ from the ground truth.
@@
Figure 12: Vorticity snapshots of the ground truth and forecasts for the EnKF and SRDA at $t=2$, 9, and 16. For comparison, the figure also shows LR simulation results without SR or DA. A low-pass filter was applied to the ground-truth snapshots to match their resolution to the LR.
@@
Figure 13 shows the time series of the mean values for the MAE ratio and MSSIM loss. Without SR or DA (green dotted lines), the test errors increase over time due to the sensitivity to initial conditions (Section 6.1). In contrast, this error growth is suppressed by the EnKF (orange dashed lines) and the SRDA (blue solid lines). The error reduction occurs every time unit since the assimilation interval was set to 1. When smoothed over time, the time series of the EnKF and SRDA have two peaks, corresponding to the two developments of small-scale patterns during the jet collapse ($t\sim 2$) and the vortex merging ($t\sim 9$). The bottom panels in Fig. 13 show the differences in the test errors between the EnKF and SRDA, where the gray areas represent the 10th to 90th percentile values. The test errors of the SRDA are significantly smaller than those of the EnKF. Thus, we conclude that the SRDA forecast outperforms the EnKF in terms of pixel-wise accuracy and pattern consistency.
@@
Figure 13: Time series of the MAE ratio and MSSIM loss for the forecast states in the 500 test simulations. The top panels show the mean test errors. The bottom panels show the differences between the SRDA and EnKF, where the gray areas cover the 10th to 90th percentile values and the solid lines indicate the mean differences.
@@
In the proposed SRDA, the LR forecast can be super-resolved to the HR at an arbitrary time using the SR model (Section 4.6). Figure 14 shows a typical vortex merging between $t=8.5$ and 11.0 from the ground truth, the LR forecast, and the super-resolved forecast. The merging of two vortices is not obvious in the LR, but is well captured in the super-resolved forecast.
@@
Figure 14: Typical vortex merging in the ground truth, the LR forecasts, and the super-resolved forecasts.
6.4 Accuracy of HR analysis
@@
Figure 15 shows snapshots of the ground truth and analyses. During the jet collapse ($t=2$), the filament structure appears much more distinct in the SRDA analysis than in the EnKF. Furthermore, during the vortex merging ($t=9$ and 11), the SRDA infers sharper vortex shapes and clearer internal structures than the EnKF.
@@
Figure 15: Vorticity snapshots of the ground truth and analyses for the EnKF and SRDA at $t=2$, 9, and 11.
@@
Figure 16 shows the time series of the MAE ratio and MSSIM loss, as in Fig. 13. The dots represent the assimilation time steps, with assimilation intervals at 1. The EnKF and SRDA analyses are inferred only at these times. The EnKF time series (orange dashed lines) show two peaks, corresponding to the two developments of small-scale patterns during the jet collapse ($t\sim 2$) and the vortex merging ($t\sim 9$). These peaks are less apparent in the SRDA time series (blue solid lines). The bottom panels in Fig. 16 indicate that the test errors of the SRDA are significantly smaller than those of the EnKF. Thus, we conclude that the SRDA analysis is more accurate than the EnKF analysis.
@@
Figure 16: Time series of the MAE ratio and MSSIM loss for the analysis states in the 500 test simulations. These states were inferred only at the analysis time steps, denoted by dots. The top panels show the mean test errors. The bottom panels show the differences between the SRDA and EnKF, where the gray areas cover the 10th to 90th percentile values and the solid lines indicate the mean differences.
@@
We compare uncertainties estimated by the EnKF and SRDA, using standard deviation (SD) as an indicator. For the EnKF, SD at each grid point was calculated from all ensemble members (i.e., 300 members). For the SRDA, SD was directly calculated by taking the square root of the estimated variances in Eq. (12).
@@
Figure 17 compares the analyses and their SDs for the EnKF and SRDA. The SRDA SD tends to be large around vortex filaments or vortex edges. This tendency appears to be most evident during the jet collapse ($t=2$). This result is reasonable because vortex locations are generally uncertain, and small changes in their locations lead to large differences in vorticity. The EnKF SD also exhibits a similar tendency; however, the SD distribution is blurred, likely reflecting the blurred vortex patterns. The SD value ranges also differ: the EnKF SD is between 0.3 and 0.9, while the SRDA SD is between 0.8 and 0.9. Compared to the EnKF, the SRDA tends to estimate larger SD in regions where vorticity is close to zero, especially at $t=9$ and 11 in Fig. 17. Although the true value of SD is unknown, our results suggest that the SRDA tends to estimate larger SD in low-vorticity regions. This tendency may be due to a simplified variance estimation in the present CVAE (Section 5.3), where the variances are computed using the SR model from the LR analysis. The SRDA demonstrated a potential for uncertainty quantification without relying on ensemble calculation; however, further model development is necessary. Recent studies demonstrate that VAEs can learn error distributions of model forecasts [22, 23, 24, 25]. Their model architectures may be effective for the SRDA.
@@
Figure 17: Vorticity snapshots of the analyses and their standard deviations (SDs) for the EnKF and SRDA at $t=2$, 9 and 11. The analyses shown here are the same as in Fig. 15.
6.5 Computation time of SRDA